# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 6.2 The compact lattice Abelian Higgs model with one-component scalar matter
The deconfinement transition line for $Q\geq 2$ is expected to connect the transition points of the models obtained for $J\to\infty$ and $\kappa\to\infty$ [41, 63]. As already mentioned, for $J\to\infty$ we obtain the ${\mathbb{Z}}_{Q}$ gauge model with Wilson action discussed in Sec. 3.5.2. For $\kappa\to\infty$, we can set $\lambda_{{\bm{x}},\mu}=1$ on all links, and thus we obtain the standard XY model. Given that the one-component model with charge $Q$ is equivalent to the gauge model with Hamiltonian  (66), which is a ${\mathbb{Z}}_{Q}$ gauge-invariant model, the finite-$J$ deconfinement transitions are expected to belong to the ${\mathbb{Z}}_{Q}$ gauge universality class  [92] (see, however, Ref. [63] for an alternative picture based on a line of FPs, thus leading to continuously varying critical exponents). As a consequence (see Sec. 3.5.2 for a discussion of the critical behavior of ${\mathbb{Z}}_{Q}$ gauge-invariant models), the continuous transitions must belong to the gauge Ising universality class and to the gauge XY universality class for $Q=2$ and $Q\geq 5$, respectively. For $Q=3$ we expect first-order transitions. For $Q=4$ the ${\mathbb{Z}}_{4}$ gauge theory that is obtained in the limit $J\to\infty$ has a gauge Ising transition, but, as discussed in Sec. 3.5.2, this is not the generic behavior expected for ${\mathbb{Z}}_{4}$ gauge invariant models. Generic ${\mathbb{Z}}_{4}$ gauge models should have continuous transitions in the gauge XY universality class. For finite $J$ we therefore expect gauge XY transitions also for $Q=4$.
@@
Figure 9: (From Ref. [92]) (Left) Plot of the third energy cumulant $C_{3}$ (its expected finite-size behavior is discussed in A.3.4) as a function of $(\kappa-\kappa_{c})L^{1/\nu_{I}}$ for $N=1$ and $Q=2$, at fixed $J=0.7$. Here $\kappa_{c}=0.5880$ and $\nu_{I}$ is the Ising value $\nu_{I}=0.629971$ (see B for a collection of estimates of $\nu_{I}$). The continuous curve is (up to nonuniversal rescaling constants) the scaling curve for $C_{3}$ in the $\mathbb{Z}_{2}$ gauge model (see Ref. [92] for a parametrization). Deviations can be explained by the presence of scaling corrections that vanish as $L^{-\omega_{I}}$ for large $L$, where $\omega_{I}=0.83$ is the universal leading scaling-correction exponent for Ising systems. This is confirmed by the plot shown in the inset, where the combination $L^{\omega_{I}}\left[L^{3-3/\nu_{I}}C_{3}(J,K,L)-{\cal C}_{3}(X)\right]$ is reported as a function of $(\kappa-\kappa_{c})L^{1/\nu_{I}}$, where ${\cal C}_{3}(X)$ (solid curve) is the asymptotic curve computed in the $\mathbb{Z}_{2}$ gauge model. The collapse of the data onto a single curve indicates that the observed scaling corrections are consistent with those expected at an Ising transition. (Right) Plot of the third energy cumulant $C_{3}$ as a function of $(\kappa-\kappa_{c})L^{1/\nu_{\rm XY}}$, for the model with $N=1$ and $Q=4$, at fixed $J=1$. Here $\kappa_{c}=1.0205$ and $\nu_{\rm XY}$ is the XY value $\nu_{\rm XY}=0.6717$ (see B for estimates of $\nu_{\rm XY}$). The continuous curve is (up to nonuniversal rescaling constants) the scaling curve for $C_{3}$ in the IXY gauge model (see Ref. [92] for a parametrization).
@@ 6.3.2 Phase diagrams for $Q\geq 2$
As shown in the right panel of Fig. 10, for small $J$ and any $\kappa\geq 0$, there is a phase in which scalar-field correlations are disordered and single-charge modes are confined (the Wilson loop obeys the area law). For large values of $J$ (low-temperature region) scalar correlations are ordered and the SU($N$) symmetry is broken. Two different phases occur here: for small $\kappa$, static single-charge test particles are confined, while they are deconfined for large $\kappa$.
@@
The transitions along the DC-OC line are analogous to those occurring along the unique transition line present for $Q=1$, see Sec. 6.3.1, and along the CM line in noncompact models. In particular, they have the same nature as in the 3D lattice CPN-1 model obtained for $\kappa=0$. Therefore, continuous transitions are only expected for $N=2$, with critical behaviors belonging to the O(3) vector universality class.
@@
In the limit $J=\infty$, the model (64) becomes equivalent to a ${\mathbb{Z}}_{Q}$ gauge model with Wilson action [86], as it occurs for $N=1$, see Sec. 6.2. A natural hypothesis is that the transitions along the OC-OD line belong to the universality class of the ${\mathbb{Z}}_{Q}$ gauge model, as in the $N=1$ compact LAH model, see Sec. 6.2. This prediction has been verified numerically for $Q=2$. Ref. [86] computed the energy cumulants, finding that they have the same FSS behavior as in the ${\mathbb{Z}}_{2}$ gauge model. These results confirm that the OC-OD transitions belong to the Ising gauge universality class for this value of $Q$.
@@
Figure 11: (Adapted from Ref. [90]) (Left) Plot of $U_{Q}$ versus $\xi_{Q}/L$, where $U_{Q}$ and $\xi_{Q}$ are the Binder parameter and the second-moment correlation length defined in terms of the gauge-invariant operator $Q_{\bm{x}}^{ab}$ defined in Eq. (57). Results for the compact and noncompact model with $N=25$ and (compact model) $Q=2$. Data are obtained in MC simulations varying $J$ across the DC-OD (compact) and CH (noncompact) line for $\kappa=1$. The solid line is an extrapolation of the data for the noncompact LAH model. Since $U_{Q}$ and $\xi_{Q}/L$ are RG invariant quantities, the collapse of the data signals that the transitions in the two models belong to the same universality class (see A.3.3 for a discussion of the FSS behavior). (Right) Same plot as in the left panel, fixing $\kappa=2$ and using compact model with $Q=3$.
@@
The transitions along the large-$\kappa$ DC-OD line have been investigated numerically in Ref. [86], which focused on the compact model for $N=15$ and $25$ and two charge values, $Q=2$ and $Q=3$, and analyzed the finite-size behavior of the correlation function of the operator $Q_{\bm{x}}^{ab}$ (some data are reported in Fig. 11). The analysis shows that the continuous DC-OD transitions belong to the same universality as the transitions along the CH line in the noncompact models discussed in Sec. 5.3.3, for all values of $N$ and $Q$ investigated. Some results taken from Ref. [86] are reported in Fig. 11, which shows that the data for the compact and the noncompact model have the same FSS behavior. Moreover, the estimated critical exponents computed in compact an noncompact models for the same value of $N$ are consistent within errors, see Fig. 7. Therefore, the stable charged FP point of the RG flow of the AHFT, see Sec. 4.1.2, which exists for $N\gtrsim 7$, controls the continuous transitions in multicomponent LAH models with both noncompact and compact gauge variables for large values of $\kappa$; more precisely, along the CH line in noncompact models and along the DC-OD line in compact models with any charge $Q$ satisfying $Q\geq 2$.
@@ 6.3.3 Relation between the compact and the noncompact lattice Abelian Higgs models
The previous results show that noncompact and compact ($Q\geq 2$) LAH models have similar phase diagrams. For $N\geq 2$, the compact DC-OC line and the noncompact CM line have the same nature, and so have the compact DC-OD and noncompact CH line for any $Q\geq 2$. The MH line and the OC-OD line both correspond to topological transitions. These results may seem unexpected. Therefore, it is worth discussing the relation between the compact and noncompact models in more detail, to explain the similarities of their phase diagram.
@@
For $Q\to\infty$, the gauge real variables $A_{{\bm{x}},\mu}$ become unbounded, and the Hamiltonian becomes equivalent to that of the noncompact formulation with $\kappa_{\rm nc}=2\kappa/Q^{2}$, where $\kappa_{\rm nc}$ is the gauge coupling of the resulting noncompact gauge-field Hamiltonian. Note that the equivalence trivially holds as long as the fluctuations of $A_{{\bm{x}},\mu}$ on each plaquette are bounded and uncorrelated for $Q\to\infty$, i.e., for any point of the phase diagram. In particular, in Sec. 3.5.2 it has already been discussed for the ${\mathbb{Z}}_{Q}$ gauge model (the limiting model of the compact LAH model for $J\to\infty$) that is equivalent to the IXY model (the limiting model of the noncompact LAH model for $J\to\infty$) in the limit considered above.
The argument presented above shows that the compact model converges to the noncompact one as $Q,\kappa\to\infty$, at fixed $\kappa_{\rm nc}=2\kappa/Q^{2}$. Therefore, also their phase diagrams must become identical for large $Q$. As far as the nature of the critical lines, the numerical evidence reported in Sec. 6.3.2 shows that the compactification of the model is an irrelevant perturbation of the noncompact model obtained for $Q=\infty$, as already discussed for the IXY model in Sec. 3.5.2. Thus, the critical behaviors become identical for sufficiently large finite $Q$. In practice, it is not really required to take large values of $Q$ to observe an IXY critical behavior. As we have already mentioned, the behavior along the DC-OC and DC-OD is the same as that on the corresponding CM and CH lines for any $Q\geq 2$, while the behavior along the OC-OD and MH lines is the same for any $Q\geq 4$. Thus, differences only occur along the line of topological transitions that end at $J=\infty$ for $Q=2$ and 3. While in the noncompact model these topological transitions belong to the IXY universality class, in the compact one IXY behavior is only observed for $Q\geq 4$. For $Q=3$ transitions are of first order, while for $Q=2$ they belong to the Ising gauge universality class.
@@ 6.3.4 Relevance of monopole configurations
In compact gauge models, the nontrivial topology of the gauge fields allows one to define topological objects. Their role has been investigated in several studies, see, e.g., Refs. [298, 213, 299, 66, 85, 89], which indicate that their presence may significantly affect the nature of the phase transitions and their critical behaviors. Refs. [85, 89] performed a numerical study to understand the role of monopole configurations, which naturally emerge in LAH models with compact gauge fields, using the monopole definition of Ref. [300].
In lattice systems with periodic boundary conditions one can define monopoles and antimopoles using the prescription proposed in Ref. [300]. This prescription starts from the noncompact lattice curl $\Theta_{{\bm{x}},\mu\nu}=\theta_{{\bm{x}},\mu}+\theta_{{\bm{x}}+\hat{\mu},\nu}-\theta_{{\bm{x}},\nu}-\theta_{{\bm{x}}+\hat{\nu},\mu}$ associated with each plaquette, where $\theta_{{\bm{x}},\mu}\in[-\pi,\pi)$ is the phase associated with compact lattice variable $\lambda_{{\bm{x}},\mu}$, $\lambda_{{\bm{x}},\mu}=e^{i\theta_{{\bm{x}},\mu}}$. Note that $\Theta_{{\bm{x}},\mu\nu}$ is antisymmetric in $\mu$ and $\nu$, so this definition provides two different quantities that differ by a sign for each plaquette, depending on the orientation. One can easily verify that, for any closed lattice surface $S$ made of elementary plaquettes, the naively defined outgoing magnetic flux across $S$ vanishes, i.e., $\sum_{P\in S}\Theta_{{\bm{x}},\mu\nu}=0$, where the sum extends to all plaquettes in $S$, $\Theta_{{\bm{x}},\mu\nu}$ is associated with plaquette $P=({\bm{x}},\mu\nu)$, and the plaquette orientation is chosen in such a way that the unit vector $\hat{\mu}\times\hat{\nu}$ points outward with respect to the surface. [^27] These rules are a discrete version of the integration rules of a differential two-form on a two-dimensional surface, which can be formalized using the formalism of lattice differential forms, see, e.g., Ref. [301]. To define a nontrivial net number of monopoles (i.e., the difference between the number of monopoles and of antimonopoles) within the surface $S$, one must isolate the singular contributions from the smooth background corresponding to small values of $|\Theta_{{\bm{x}},\mu\nu}|$. This can be done by defining
@@
A topic that has been investigated in the literature is whether the suppression of the monopole configurations in the partition function changes the nature of the transitions, or whether it gives rise to new universality classes somehow related to those found in noncompact formulations, in which analogous monopole configurations are absent.
@@
These issues have been addressed in Refs. [85, 89]. In the lattice CPN-1 model defined by the Hamiltonian reported in Eq. (8) a finite density of monopoles is observed in the disordered low-$J$ phase up to the critical point, while in the low-temperature regime $J>J_{c}$ only isolated pairs of monopoles and anti-monopoles are typically present, whose number decreases rapidly when increasing $J$, since $\theta_{{\bm{x}},\mu}$ approaches zero in the large-$J$ limit. Thus, the monopole density is an appropriate order parameter of the transition. As a consequence, by considering only monopole-free configurations, one is changing the nature of the small-$J$ phase, and therefore one expects a different critical behavior or even the absence of any transition.
@@
To define a monopole-free version of the CPN-1 model (MFCPN-1), one may change the configuration space, considering only configurations for which $N_{\rm mono}(C)=0$ on any elementary lattice cube $C$ [85, 89]. The numerical analyses of Refs. [85, 89] show that the MFCPN-1 models, and some extensions obtained by relaxing the unit-length condition on the scalar fields, undergo finite-$J$ transitions. However, the features of these transitions are apparently unrelated with those of the transitions in the standard CPN-1 model. In particular, there is no O(3) continuous transition for $N=2$. Moreover, the numerical data definitely exclude that the MFCPN-1 model has a transition associated with the AHFT for large values of $N$  [85, 89].
@@
Ref. [86] studied the critical behavior of monopole-free “higher-charge” CPN-1 models, obtained by setting $\kappa=0$ in Eq. (64). [^28] If we set $\kappa=0$ in Eq. (64) we obtain a charge-$Q$ CPN-1 model. In the absence of the monopole-free condition, this model is equivalent to the standard $Q=1$ CPN-1 model (it is enough to redefine the gauge fields). This equivalence does not hold in the monopole-free version, as in this case it it is not possible to replace $\lambda_{{\bm{x}},\mu}^{Q}$ with $\lambda_{{\bm{x}},\mu}$, as the gauge fields also appear in the monopole-free condition. Thus, we obtain a different monopole-free model for each $Q$. The results show that, for $Q\geq 2$ the monopole-free model has the same behavior as the standard CPN-1 model. This means that in the compact LAH model the topological properties of the gauge field are inessential for any $Q\geq 2$, which is consistent with the fact that the compact and noncompact model have analogous phase diagrams for $Q\geq 2$, and in particular with the appearance (for $N\gtrsim 7$) of an AHFT critical behavior both in the noncompact LAH model along the CH line and in the compact LAH model with $Q\geq 2$ along the DC-OD line, see the discussion in Sec. 6.3.2.

### v3

@@ 6.2 The compact lattice Abelian Higgs model with one-component scalar fields
The deconfinement transition line for $Q\geq 2$ is expected to connect the transition points of the models obtained for $J\to\infty$ and $\kappa\to\infty$ [41, 63]. As already mentioned, for $J\to\infty$ we obtain the ${\mathbb{Z}}_{Q}$ gauge model with Wilson action discussed in Sec. 3.5.2. For $\kappa\to\infty$, we can set $\lambda_{{\bm{x}},\mu}=1$ on all links, and thus we obtain the standard XY model. Given that the one-component model with charge $Q$ is equivalent to the gauge model with Hamiltonian  (69), which is a ${\mathbb{Z}}_{Q}$ gauge-invariant model, the finite-$J$ deconfinement transitions are expected to belong to the ${\mathbb{Z}}_{Q}$ gauge universality class  [92] (see, however, Ref. [63] for an alternative picture based on a line of FPs, thus leading to continuously varying critical exponents). As a consequence (see Sec. 3.5.2 for a discussion of the critical behavior of ${\mathbb{Z}}_{Q}$ gauge-invariant models), the continuous transitions must belong to the Ising gauge universality class and to the XY gauge universality class for $Q=2$ and $Q\geq 5$, respectively. For $Q=3$ we expect first-order transitions. For $Q=4$ the ${\mathbb{Z}}_{4}$ gauge theory that is obtained in the limit $J\to\infty$ has an Ising gauge transition, but, as discussed in Sec. 3.5.2, this is not the generic behavior expected for ${\mathbb{Z}}_{4}$ gauge invariant models. Generic ${\mathbb{Z}}_{4}$ gauge models should have continuous transitions in the XY gauge universality class. For finite $J$ we therefore expect XY gauge transitions also for $Q=4$.
@@
Figure 9: (From Ref. [92]) (Left) Plot of the third energy cumulant $C_{3}$ (its expected finite-size behavior is discussed in A.3.2) as a function of $(\kappa-\kappa_{c})L^{1/\nu_{I}}$ for $N=1$ and $Q=2$, at fixed $J=0.7$. Here $\kappa_{c}=0.5880$ and $\nu_{I}$ is the Ising value $\nu_{I}=0.629971$ (see B for a collection of estimates of $\nu_{I}$). The continuous curve is (up to nonuniversal rescaling constants) the scaling curve for $C_{3}$ in the $\mathbb{Z}_{2}$ gauge model (see Ref. [92] for a parametrization). Deviations can be explained by the presence of scaling corrections that vanish as $L^{-\omega_{I}}$ for large $L$, where $\omega_{I}=0.83$ is the universal leading scaling-correction exponent for Ising systems. This is confirmed by the plot shown in the inset, where the combination $L^{\omega_{I}}\left[L^{3-3/\nu_{I}}C_{3}(J,K,L)-{\cal C}_{3}(X)\right]$ is reported as a function of $(\kappa-\kappa_{c})L^{1/\nu_{I}}$, where ${\cal C}_{3}(X)$ (solid curve) is the properly rescaled asymptotic curve computed in the $\mathbb{Z}_{2}$ gauge model. The collapse of the data onto a single curve indicates that the observed scaling corrections are consistent with those expected at an Ising transition. (Right) Plot of the third energy cumulant $C_{3}$ as a function of $(\kappa-\kappa_{c})L^{1/\nu_{\rm XY}}$, for the model with $N=1$ and $Q=4$, at fixed $J=1$. Here $\kappa_{c}=1.0205$ and $\nu_{\rm XY}$ is the XY value $\nu_{\rm XY}=0.6717$ (see B for estimates of $\nu_{\rm XY}$). The continuous curve is (up to nonuniversal rescaling constants) the scaling curve for $C_{3}$ in the IXY gauge model (see Ref. [92] for a parametrization).
@@ 6.3.2 Phase diagram for $Q\geq 2$
As shown in the right panel of Fig. 10, for small $J$ and any $\kappa\geq 0$, there is a phase in which scalar-field correlations are disordered and single-charge modes are confined (the Wilson loop obeys the area law). For large values of $J$ (low-temperature region) scalar correlations are ordered and the SU($N$) symmetry is broken. Two different phases occur here: static single-charge test particles are confined for small $\kappa$ and deconfined for large $\kappa$.
@@
The transitions along the DC-OC line are analogous to those occurring along the unique transition line present for $Q=1$, see Sec. 6.3.1, and along the CM line in noncompact models. In particular, they have the same nature as in the 3D lattice CPN-1 model obtained for $\kappa=0$. Therefore, continuous transitions are only expected for $N=2$, with a critical behavior belonging to the O(3) vector universality class.
@@
In the limit $J\to\infty$, the model (67) becomes equivalent to a ${\mathbb{Z}}_{Q}$ gauge model with Wilson action [86], as it occurs for $N=1$, see Sec. 6.2. A natural hypothesis is that the transitions along the OC-OD line belong to the universality class of the ${\mathbb{Z}}_{Q}$ gauge model, as in the $N=1$ compact LAH model, see Sec. 6.2. This prediction has been verified numerically for $Q=2$. Ref. [86] computed the energy cumulants, finding that they have the same FSS behavior as in the ${\mathbb{Z}}_{2}$ gauge model. These results confirm that the OC-OD transitions belong to the Ising gauge universality class for this value of $Q$.
@@
Figure 11: (Adapted from Ref. [90]) (Left) Plot of $U_{Q}$ versus $\xi_{Q}/L$, where $U_{Q}$ and $\xi_{Q}$ are the Binder parameter and the second-moment correlation length defined in terms of the gauge-invariant operator $Q_{\bm{x}}^{ab}$ defined in Eq. (60). Results for the compact and noncompact model with $N=25$ and (compact model) $Q=2$. Data are obtained in MC simulations varying $J$ across the DC-OD (compact) and CH (noncompact) line for $\kappa=1$. The solid line is an extrapolation of the data for the noncompact LAH model. Since $U_{Q}$ and $\xi_{Q}/L$ are RG invariant quantities, the collapse of the data signals that the transitions in the two models belong to the same universality class (see A.3.4 for a discussion of the FSS behavior). (Right) Same plot as in the left panel for the compact model with $Q=3$, fixing $\kappa=2$.
@@
The transitions along the large-$\kappa$ DC-OD line have been investigated numerically in Ref. [86], which focused on the compact model for $N=15$ and $25$ and two charge values, $Q=2$ and $Q=3$, and analyzed the finite-size behavior of the correlation function of the operator $Q_{\bm{x}}^{ab}$ (some data are reported in Fig. 11). The analysis shows that the continuous DC-OD transitions belong to the same universality as the transitions along the CH line in the noncompact models discussed in Sec. 5.3.3, for all values of $N$ and $Q$ investigated. Some results taken from Ref. [86] are reported in Fig. 11, which shows that the data for the compact and the noncompact model have the same FSS behavior. Moreover, the estimated critical exponents computed in the compact and in the noncompact model for the same value of $N$ are consistent within errors, see Fig. 7. Therefore, the stable charged FP point of the RG flow of the AHFT, see Sec. 4.1.2, which exists for $N\gtrsim 7$, controls the continuous transitions in multicomponent LAH models with both noncompact and compact gauge variables for large values of $\kappa$; more precisely, along the CH line in noncompact models and along the DC-OD line in compact models with any charge $Q$ satisfying $Q\geq 2$.
@@ 6.3.3 Relation between the compact and the noncompact lattice Abelian Higgs models
The previous results show that noncompact and compact ($Q\geq 2$) LAH models have similar phase diagrams. For $N\geq 2$, the compact DC-OC line and the noncompact CM line have the same nature, and so have the compact DC-OD and noncompact CH line for any $Q\geq 2$. The MH line and the OC-OD line both correspond to topological transitions. These results may appear unexpected. Therefore, it is worth discussing the relation between the compact and noncompact models in more detail, to explain the similarities of their phase diagram.
@@
For $Q\to\infty$, the gauge real variables $A_{{\bm{x}},\mu}$ become unbounded, and the Hamiltonian becomes equivalent to that of the noncompact formulation with $\kappa_{\rm nc}=2\kappa/Q^{2}$, where $\kappa_{\rm nc}$ is the gauge coupling of the resulting noncompact gauge-field Hamiltonian. Note that this equivalence trivially holds as long as the fluctuations of $A_{{\bm{x}},\mu}$ on each plaquette are bounded and uncorrelated for $Q\to\infty$, i.e., for any point of the phase diagram. In particular, in Sec. 3.5.2 it has already been discussed for the ${\mathbb{Z}}_{Q}$ gauge model (the limiting model of the compact LAH model for $J\to\infty$), which is equivalent to the IXY model (the limiting model of the noncompact LAH model for $J\to\infty$) in the limit considered above.
The argument presented above shows that the compact model converges to the noncompact one as $Q,\kappa\to\infty$, at fixed $\kappa_{\rm nc}=2\kappa/Q^{2}$. Therefore, also their phase diagrams must become identical for large $Q$. As far as the nature of the critical lines, the numerical evidence reported in Sec. 6.3.2 shows that the compactification of the model is an irrelevant perturbation of the noncompact model obtained for $Q=\infty$, as already discussed for the IXY model in Sec. 3.5.2. Thus, the critical behavior should become identical for $Q\to\infty$. In practice, it is not really required to take large values of $Q$ to observe the same IXY critical behavior, as in the noncompact model. As we have already mentioned, the behavior along the DC-OC and DC-OD is the same as that on the corresponding CM and CH lines for any $Q\geq 2$, while the behavior along the OC-OD and MH lines is the same for any $Q\geq 4$. Thus, differences only occur along the line of topological transitions that end at $J=\infty$ for $Q=2$ and 3. While in the noncompact model these topological transitions belong to the IXY universality class, in the compact one the IXY critical behavior is only observed for $Q\geq 4$. For $Q=3$ transitions are of first order, while for $Q=2$ they belong to the Ising gauge universality class.
@@ 6.3.4 Relevance of monopole configurations
In compact gauge models, the nontrivial topology of the gauge fields allows one to define topological objects. Their role has been investigated in several studies, see, e.g., Refs. [326, 242, 327, 66, 85, 89], which indicate that their presence may significantly affect the nature of the phase transitions and their critical behavior. Refs. [85, 89] performed a numerical study to understand the role of monopole configurations, which naturally emerge in LAH models with compact gauge fields, using the monopole definition of Ref. [328].
In lattice systems with periodic boundary conditions one can define monopoles and antimopoles using the prescription proposed in Ref. [328]. This prescription starts from the noncompact lattice curl $\Theta_{{\bm{x}},\mu\nu}=\theta_{{\bm{x}},\mu}+\theta_{{\bm{x}}+\hat{\mu},\nu}-\theta_{{\bm{x}},\nu}-\theta_{{\bm{x}}+\hat{\nu},\mu}$ associated with each plaquette, where $\theta_{{\bm{x}},\mu}\in[-\pi,\pi)$ is the phase associated with the compact lattice variable $\lambda_{{\bm{x}},\mu}$, $\lambda_{{\bm{x}},\mu}=e^{i\theta_{{\bm{x}},\mu}}$. Note that $\Theta_{{\bm{x}},\mu\nu}$ is antisymmetric in $\mu$ and $\nu$, so this definition provides two different quantities that differ by a sign for each plaquette, depending on the orientation. One can easily verify that, for any closed lattice surface $S$ made of elementary plaquettes, the naively defined outgoing magnetic flux across $S$ vanishes, i.e., $\sum_{P\in S}\Theta_{{\bm{x}},\mu\nu}=0$, where the sum extends to all plaquettes in $S$, $\Theta_{{\bm{x}},\mu\nu}$ is associated with plaquette $P=({\bm{x}},\mu\nu)$, and the plaquette orientation is chosen in such a way that the unit vector $\hat{\mu}\times\hat{\nu}$ points outward with respect to the surface. [^28] These rules are a discrete version of the integration rules of a differential two-form on a two-dimensional surface, which can be formalized using the formalism of lattice differential forms, see, e.g., Ref. [329]. To define a nontrivial net number of monopoles (i.e., the difference between the number of monopoles and of antimonopoles) within the surface $S$, one must isolate the singular contribution from the smooth background corresponding to small values of $|\Theta_{{\bm{x}},\mu\nu}|$. This can be done by defining
@@
A topic that has been investigated in the literature is whether the suppression of the monopole configurations in the partition function changes the nature of the transition, or whether it gives rise to new universality classes somehow related to those found in noncompact formulations, in which analogous monopole configurations are absent. These issues have been addressed in Refs. [85, 89]. In the lattice CPN-1 model defined by the Hamiltonian reported in Eq. (8), a finite density of monopoles is observed in the disordered low-$J$ phase up to the critical point, while in the low-temperature regime $J>J_{c}$ only isolated pairs of monopoles and anti-monopoles are typically present, whose number decreases rapidly with increasing $J$, since $\Theta_{{\bm{x}},\mu\nu}$ approaches zero in the large-$J$ limit. Thus, the monopole density is an appropriate order parameter of the transition. As a consequence, by considering only monopole-free configurations, one is changing the nature of the small-$J$ phase, and therefore one expects a different critical behavior or even the absence of any transition.
@@
To define a monopole-free version of the CPN-1 model (MFCPN-1), one may change the configuration space, considering only configurations for which $N_{\rm mono}(C)=0$ on any elementary lattice cube $C$ [85, 89]. The numerical analyses of Refs. [85, 89] show that the MFCPN-1 models, as well as some extensions obtained by relaxing the unit-length condition on the scalar fields, undergo finite-$J$ transitions. However, the features of these transitions are apparently unrelated with those of the transitions in the standard CPN-1 model. In particular, there is no O(3) continuous transition for $N=2$. Moreover, the numerical data definitely exclude that the MFCPN-1 model has a transition associated with the AHFT for large values of $N$  [85, 89].
@@
Ref. [86] also studied the critical behavior of monopole-free “higher-charge” CPN-1 models, obtained by setting $\kappa=0$ in Eq. (67). [^29] If we set $\kappa=0$ in Eq. (67) we obtain a charge-$Q$ CPN-1 model. In the absence of the monopole-free condition, this model is equivalent to the standard $Q=1$ CPN-1 model (it is enough to redefine the gauge fields). This equivalence does not hold in the monopole-free version, as in this case it is not possible to replace $\lambda_{{\bm{x}},\mu}^{Q}$ with $\lambda_{{\bm{x}},\mu}$, as the gauge fields also appear in the monopole-free condition. Thus, we obtain a different monopole-free model for each $Q$. The results show that, for $Q\geq 2$ the monopole-free model has the same behavior as the standard CPN-1 model. This means that in the compact LAH model the topological properties of the gauge field are inessential for any $Q\geq 2$, which is consistent with the fact that the compact and noncompact model have analogous phase diagrams for $Q\geq 2$, and in particular with the appearance (for $N\gtrsim 7$) of an AHFT critical behavior both in the noncompact LAH model along the CH line and in the compact LAH model with $Q\geq 2$ along the DC-OD line, see the discussion in Sec. 6.3.2.